at a bicycle store there were a bunch of bicycles and tricycles. If there were 32 seats and 72 wheels, how many bicycles and how many tricycles were there
Answer by edjones(8007)

Let t be the number of tricycles.
tricycle wheels + bicycle wheels = 72
3t+2(32-t)=72
3t+64-2t=72
t=8
32-8=24 bicycles
.
Ed